In a normal distribution 10% of the items are under 50 and 90% of the items are under 

80. Find the values of mean and standard deviation of the distribution.


1
Expert's answer
2021-07-01T13:18:16-0400

Solution:

Value of z = "\frac{x - \bar{x} }{\sigma }"

Value of z corresponding to the first group of items under 10"\%" = 0.5 – 0.1 = 0.4

0.4 on the table = - 0.6554

Z = "\frac{x - \bar{x} }{\sigma }"

-0.6554 = "\frac{50 - \bar{x} }{\sigma }"

"-0.6554\sigma = 50 - \bar{x}"

"\bar{x}-0.6554\sigma = 50"

Value of z corresponding to the second group of items under 90"\%" = 1 – 0.9 = 0.1

0.1 on the table = 0.5398

Z = "\frac{x - \bar{x} }{\sigma }"

0.5398= "\frac{90 - \bar{x} }{\sigma }"

"0.5398\sigma = 90 - \bar{x}"

"\bar{x} + 0.5398\sigma = 90"

Solve for the system equations:

"\bar{x}-0.6554\sigma = 50"

"\bar{x} + 0.5398\sigma = 90"

"-1.1952\sigma = -40"

"\sigma = 33.47"

The value of the standard deviation in the distribution = 33.47


Substitute the value of "\sigma" in the first equation to derive x̅:

"\bar{x}-0.6554\sigma = 50"

"\bar{x}-0.6554(33.47) = 50"

"\bar{x} - 21.94 = 50"

"\bar{x}= 50 + 21.94 = 71.94"

"\bar{x} = 71.94"

The value of the mean of the distribution = 71.94